OpenStudy (anonymous):

Want a fan and medal? Help me out! If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (hanner_b_nanner):

You just plug in 3 wherever you see x!!! g(3)=2(3)+2 g(3)=6+2 x=8

OpenStudy (anonymous):

Thanks so much beautiful :)

OpenStudy (hanner_b_nanner):

anytimeee! :D